Please can you help me with the following problem.
Use the geometric sequence of numbers 1, 2, 4, 8,…to find the following:
a) What is r, the ratio between 2 consecutive terms?
b) Using the formula for the nth term of a geometric sequence, what is the 24th term?
c) Using the formula for the sum of a geometric series, what is the sum of the first 10 terms?
Thank you for your help.

Answer by Nate(3500)

1, 2, 4, 8....
r = a selection / the previous number
r = 2/1 = 4/2 = 8/4
r = 2
...
(2)^(n - 1) where n = 24
2^(23) = 8,388,608
...
a(1 - r^n)/(1 - r)
(1 - 2^n)/(1 - 2)
2^10 - 1
1023